# QoE Maximization for Multiple-UAV-Assisted Multi-Access Edge Computing: An Online Joint Optimization Approach

### v1

@@ Abstract
In disaster scenarios, conventional terrestrial multi-access edge computing (MEC) paradigms, which rely on fixed infrastructure, may become unavailable due to infrastructure damage. With high-probability line-of-sight (LoS) communication, flexible mobility, and low cost, unmanned aerial vehicle (UAV)-assisted MEC is emerging as a new promising paradigm to provide edge computing services for ground user devices (UDs) in disaster-stricken areas. However, the limited battery capacity, computing resources, and spectrum resources also pose serious challenges for UAV-assisted MEC, which can potentially shorten the service time of UAVs and degrade the quality of experience (QoE) of UDs without an effective control approach. To this end, in this work, we first present a hierarchical architecture of multiple-UAV-assisted MEC networks that enables the coordinated provision of edge computing services by multiple UAVs. Then, we formulate a joint task offloading, resource allocation, and UAV trajectory planning optimization problem (JTRTOP) to maximize the QoE of UDs while considering the energy consumption constraints of UAVs. Since the problem is proven to be a future-dependent and NP-hard problem, we propose a novel online joint task offloading, resource allocation, and UAV trajectory planning approach (OJTRTA) to solve the problem. Specifically, the JTRTOP is first transformed into a per-slot real-time optimization problem (PROP) using the Lyapunov optimization framework. Then, a two-stage optimization method based on game theory and convex optimization is proposed to solve the PROP. Simulation results provide empirical evidence supporting the superior system performance of the proposed OJTRTA in comparison to alternative approaches.
@@
UAV-assisted MEC network, task offloading, resource allocation, trajectory planning, game theory.
1 Introduction
@@
With the advancements in artificial intelligence and wireless communication technologies, many intelligent applications with strict requirements on computing resources and latency have emerged rapidly, such as real-time video analysis [2], virtual and augmented realities [3], and interactive online games [4]. However, the limited battery capacity and computing capability of user devices (UDs) make it difficult to maintain high-level quality of experience (QoE) for these intelligent applications [5]. To overcome this challenge, multi-access edge computing (MEC) has emerged as a promising paradigm to provide reliable edge computing services by deploying MEC servers at the edge of the network [6]. Specifically, UDs can offload latency-sensitive and computation-hungry tasks to nearby MEC servers to improve the QoE. However, conventional terrestrial MEC heavily relies on the deployment of ground infrastructure, which can become unavailable in the event of a disaster where ground-based infrastructure may be damaged.
@@
The limitations of conventional terrestrial MEC have prompted a paradigm shift toward unmanned aerial vehicle (UAV)-assisted MEC due to the high probability line-of-sight (LoS) communication, flexible mobility, and low cost [7, 8]. Specifically, the high probability of LoS links for UAVs enhances the communication coverage, network capacity, and reliable connectivity [9, 10]. Moreover, the flexible mobility of UAVs enables rapid and on-demand deployment, which is crucial in disaster scenarios. Besides, the low cost makes the UAV-assisted MEC system feasible and scalable. Therefore, UAV-assisted MEC holds tremendous potential for providing edge computing services to UDs in disaster-stricken areas.
@@
However, there exist several fundamental challenges that need to be addressed to fully exploit the benefits of UAV-enabled MEC. i) Resource Allocation. Various tasks of UDs are generally heterogeneous and time-varying, and they have stringent requirements for the computing service. However, the limited computing and spectrum resources of UAV-enabled MEC and the stringent demands of UDs could lead to the competition for resources inside the MEC server, especially during peak times. Consequently, devising an efficient resource allocation strategy to fulfill the diverse task demands under resource constraints poses a significant challenge for the MEC server. ii) Task Offloading. Due to resource competition, the offloading decisions of each UD depends not only on its own offloading requirements but also on the offloading decisions of the other UDs, resulting in coupled and complex offloading decision-making among UDs. iii) Trajectory Planning. While the mobility of UAVs augments the flexibility and elasticity of MEC, it also presents notable challenges in terms of UAV trajectory planning. iv) Energy Constraint. The limited onboard battery capacity of UAVs leads to finite service time, which makes it challenging to balance the service time of UAVs and QoE of UDs. In addition, within the resource and energy limitations of UAVs, the interdependencies among resource allocation decisions, task offloading decisions, and trajectory planning introduce intricate coupling relationships that contribute to the complexity of the decision-making process.
To address the challenges described above, we present a novel online approach that integrates task offloading, resource allocation, and UAV trajectory planning for the purpose of maximizing the QoE of UDs while adhering to the energy consumption constraint of UAVs. The key contributions of this work are outlined as follows:
@@
  • System Architecture. We propose a hierarchical architecture for the multiple-UAV-assisted MEC network to efficiently coordinate multiple UAVs in providing aerial edge computing services for UDs. Furthermore, we take into account the dynamic mobility and time-varying computational demands of UDs to accurately capture the features of real-world edge computing scenarios within the proposed system.
@@
  • Problem Formulation. We formulate a novel joint task offloading, resource allocation, and UAV trajectory planning optimization problem (JTRTOP), with the aim of maximizing the QoE of UDs under the UAV energy consumption constraint. Specifically, the QoE of UDs is theoretically quantified by synthesizing the task completion delay and UD energy consumption. Moreover, the optimization problem is proved to be future-dependent and NP-hard.
@@
  • Algorithm Design. To solve the formulated problem, we propose an online joint task offloading, resource allocation, and UAV trajectory planning approach (OJTRTA). Specifically, we first transform the JTRTOP into a per-slot real-time optimization problem (PROP) by using the Lyapunov optimization framework. Then, we propose a two-stage method to optimize the task offloading, resource allocation and UAV position of PROP by using convex optimization and game theory.
@@
  • Validation. The effectiveness and performance of the proposed OJTRTA are validated through both theoretical analysis and simulation experiments. In particular, the theoretical analysis establishes that the proposed OJTRTA satisfies both the energy consumption constraint of UAVs and achieves convergence to a sub-optimal solution within polynomial time. Additionally, the simulation results affirm the superiority of OJTRTA over alternative approaches.
@@
The subsequent sections of this work are structured as follows. In Section 2, an overview of the related work is provided. Section 3 details the relevant system models and problem formulation. Section 4 describes the Lyapunov-based problem transformation. Section 5 presents the two-stage optimization method and theoretical analysis. In Section 6, simulation results are displayed and analyzed. Finally, this work is concluded in Section 7.
2 Related Work
@@
In this section, we provide a comprehensive review of the relevant studies pertaining to UAV-enabled MEC architecture, formulation of joint optimization problems, and online optimization approach. Moreover, we outline the distinctions between this work and the related works in Table I.
2.1 UAV-enabled MEC Architecture
@@
UAV-enabled MEC is a promising paradigm to dynamically expand the computational capabilities of MEC networks and facilitating emergency scenarios. From the perspective of network architecture, the existing works can primarily be categorized into single UAV-enabled MEC architecture and multiple-UAV-enabled MEC architecture.
@@
For single UAV-assisted MEC networks, Lin et al. [11] explored the maximization of UAV energy efficiency while simultaneously considering fairness in task offloading within a single UAV-assisted MEC network. Yu et al. [12] presented a novel single UAV-assisted MEC network, where a UAV is deployed to provide MEC services to Internet of Things (IoT) devices in areas that are inaccessible to the edge cloud due to ground signal blockage or shadowing. However, a single UAV possesses insufficient coverage range and resources, making it challenging to provide effective services in large-scale areas or areas with a high density of users.
@@
Multiple-UAV-enabled MEC networks, which harnesses the computational and communication capabilities of multiple UAVs, has been increasingly gaining attention. For example, Li et al. [13] investigated a multiple-UAV-assisted MEC system and devised a robust approach for optimizing computation offloading and trajectory planning. Shidrokh et al. [14] conducted a study on task offloading and resource allocation in a multiple-UAV-enabled MEC network, and proposed a two-layer collaborative evolution model to reduce energy consumption and task computation time. Bai et al. [15] proposed a multiple-UAV-enabled edge-cloud computing system that leverages the collaboration between UAVs and remote clouds to deliver exceptional computational capabilities.
@@
2.2 Formulation of Joint Optimization Problems
@@ 2.2.1 Optimization Objectives
Several studies are dedicated to minimizing task completion latency. For instance, Luan et al. [16] formulated a joint optimization problem for topology reconstruction and subtask scheduling to minimize the average completion time of subtasks in a UAV-assisted MEC emergency system. Deng et al. [17] investigated the minimization of task service latency in air-ground integrated wireless networks, considering constraints on learning accuracy and energy consumption.
@@
Several studies focus on minimizing the energy consumption. For example, in the context of an multiple-UAV-assisted two-stage MEC system, Ei et al. [18] aimed to minimize the energy consumption of both mobile devices and UAVs under constraints of task tolerance latency and available resources. Liu et al. [19] conducted a study on a multi-input single-output UAV-assisted MEC network and presented a system energy minimization problem.
@@
There are also the studies that consider the combination of task completion latency and energy consumption as the optimization objective. For example, Tong et al. [20] introduced a UAV-enabled multi-hop cooperative paradigm and formulated an optimization problem to maximize the system utility. The system utility was constructed by jointly considering task completion latency and UAV energy consumption. Hao et al. [21] explored the problem of maximizing the long-term average system gain in a UAV-assisted MEC system, where the system gain was theoretically constructed by combining task completion latency and system energy consumption.
@@
This work differs from the aforementioned studies in the following aspects. Firstly, this work considers a more comprehensive set of performance metrics, including task completion latency, UD energy consumption, and UAV energy consumption, which aligns better with the practical characteristics of UAV-assisted MEC networks. Secondly, we combine task completion latency and UD energy consumption into a UD cost function as the optimization objective, while treating UAV energy consumption as a constraint. By formulating the optimization problem in this manner, we can simultaneously enhance the QoE of UDs and ensure reliable service time for UAVs.
2.2.2 Optimization Variables
@@
Researchers have extensively explored various aspects of UAV-assisted MEC systems, with particular emphasis on task offloading, resource allocation, and UAV trajectory planning. For example, Chen et al. [22] conducted research on task offloading for IoT nodes and trajectory planning for multiple UAVs to reduce UAV energy consumption. Song et al. [23] focused on investigating task offloading and UAV trajectory control problems to improve the performance of UAV-assisted MEC systems. To minimize energy consumption and latency, Pervez et al. [24] jointly optimized task offloading, CPU frequency allocation, transmission power, and UAV trajectory. Seid et al. [25] explored the joint optimization of task offloading, age of information, computation resource allocation, and communication resource allocation to enhance the performance of UAV-enabled Internet of Medical Things networks.
@@
These previous studies have several limitations. First, task offloading, resource allocation, and UAV trajectory planning are key factors in enhancing the performance of UAV-assisted MEC networks. However, the aforementioned studies did not comprehensively optimize these factors, which hinders the full exploitation of the advantages of UAV-assisted MEC. Furthermore, these studies consider resource allocation from either the communication or the computation aspects, which may lead to severe performance degradation in practical UAV-enabled MEC systems where both communication and computing resources are insufficient. Motivated by these issues, in this work, we jointly optimize task offloading, communication and computation resource allocation, as well as UAV trajectory planning.
TABLE I: Comparison between Related Works with This Work
UAV-enabled MEC
architecture
Optimization objectives Optimization variables Online approach design
Reference
Hierarchical multi-
UAV architecture
Latency
Energy
consumption
Task
offloading
Resource
allocation
Trajectory
planning
Lyapunov
optimization
Game
theory
Convex
optimization
[11]
[12]
[13]
[14]
[15]
[16]
[17]
[18]
[19]
[20]
[21]
[22]
[23]
[24]
[25]
[26]
[27]
[28]
[29]
[30]
[31]
[32]
[33]
This work
2.3 Online Optimization Approach
@@
To tackle the intricate joint optimization problems of task offloading, resource allocation, and trajectory planning for UAV-assisted MEC systems, numerous works have employed offline approaches for system design. These offline approaches formulate solutions under the assumption that the user’s location remains unchanged and the user’s demands are either fixed or known in advance [26, 27, 28]. However, in many edge computing scenarios, such as online games and real-time video analysis, computation demands arrive in a stochastic manner, and user locations exhibit dynamic changes. Therefore, it is necessary to develop online approaches for UAV-enabled MEC to make real-time decisions without knowing future information.
@@
Several studies have also explored online approaches. For example, Zhou et al. [29] developed a two time-scale online approach for caching and task offloading by leveraging the Lyapunov optimization framework. Considering the time-varying computing requirements of user equipment, Wang et al. [30] jointly optimized the user association, resource allocation and trajectory of UAVs with the aim of minimizing energy consumption of all user equipment. To minimize the average power consumption of the system with randomly arriving user tasks, Hoang et al. [31] developed a Lyapunov-guided deep reinforcement learning (DRL) framework. Miao et al. [32] proposed a deep deterministic policy gradient (DDPG)-based algorithm to optimize the computational resources allocation and UAV flight trajectory for UAV-assisted MEC. Xu et al. [33] formulated a long-term optimization problem for the joint optimization of UAV trajectory and computation resource allocation and proposed a trajectory planning algorithm based on proximal policy optimization (PPO).
@@
The Lyapunov-based optimization framework and DRL represent two viable methodologies for developing online approaches. While DRL is a powerful technique for training agents to make real-time decisions, it necessitates a substantial amount of sample data to learn optimal strategies and suffers from a lack of interpretability. In contrast, the Lyapunov-based optimization framework does not depend on sample data and provides stable performance guarantees. Therefore, we employ the Lyapunov-based optimization framework to devise our online approach. In a departure from existing research, we propose a novel two-stage optimization algorithm based on game theory and convex optimization within the Lyapunov-based optimization framework. This algorithm demonstrates both low computational complexity and superior performance.
3 System Model and Problem Formulation
Figure 1: The hierarchical multiple-UAV-assisted MEC system consists of an aerial control layer, an aerial edge layer and a ground UD layer.

### v2

# Supplemental-Material

@@ Abstract
In disaster scenarios, conventional terrestrial multi-access edge computing (MEC) paradigms, which rely on ground infrastructure, may become unavailable due to infrastructure damage. With high-probability line-of-sight (LoS) communication, flexible mobility, and low cost, unmanned aerial vehicle (UAV)-assisted MEC is emerging as a promising paradigm to provide edge computing services for ground user devices (UDs) in disaster-stricken areas. However, the limited battery capacity, computing resources, and spectrum resources also pose serious challenges for UAV-assisted MEC, which can potentially shorten the service time of UAVs and degrade the quality of experience (QoE) of UDs without an effective control approach. To this end, in this work, we first present a hierarchical architecture of multiple-UAV-assisted MEC networks that enables the coordinated provision of edge computing services by multiple UAVs. Then, we formulate a joint task offloading, resource allocation, and UAV trajectory control optimization problem (JTRTOP) to maximize the QoE of UDs while considering the energy and resource constraints of UAVs. Since the problem is proven to be a future-dependent and NP-hard problem, we propose a novel online joint task offloading, resource allocation, and UAV trajectory control approach (OJTRTA) to solve the problem. Specifically, the JTRTOP is first transformed into a per-slot real-time optimization problem (PROP) using the Lyapunov optimization framework. Then, a two-stage optimization method based on game theory and convex optimization is proposed to solve the PROP. Simulation results show that the proposed OJTRTA outperforms various benchmark approaches and achieves at least a 10% improvement in the QoE of UDs compared to deep reinforcement learning (DRL)-based algorithms, thereby validating the superiority of the proposed approach.
@@
UAV-assisted MEC, task offloading, resource allocation, UAV trajectory control, game theory.
I Introduction
@@
With the development of wireless communication technologies and artificial intelligence, many intelligent applications with strict requirements on computing resources and latency have emerged rapidly, such as real-time video analysis, virtual and augmented realities, and interactive online games [2]. However, the limited computing capabilities and battery capacity of user devices (UDs) make it difficult to maintain high-level quality of experience (QoE) for these intelligent applications [3]. To overcome this challenge, multi-access edge computing (MEC) has emerged as a promising paradigm to offer reliable edge computing services by deploying MEC servers at the network edge [4]. Specifically, UDs can offload computation-hungry and latency-sensitive tasks to nearby MEC servers for execution, which can effectively reduce task completion delay and energy consumption, thereby enhancing the QoE of UDs. However, conventional terrestrial MEC heavily relies on ground infrastructure, which can become unavailable in the event of a disaster where ground infrastructure may be damaged.
@@
The limitations of conventional terrestrial MEC have prompted a paradigm shift toward unmanned aerial vehicle (UAV)-assisted MEC due to the high probability line-of-sight (LoS) communication, flexible mobility, and low cost [5]. Specifically, the high probability of LoS links for UAVs enhances the communication coverage, network capacity, and reliable connectivity. Moreover, the flexible mobility of UAVs enables rapid and on-demand deployment, which is crucial in disaster scenarios. Finally, the low cost makes the UAV-assisted MEC system feasible and scalable. Therefore, UAV-assisted MEC holds tremendous potential for providing edge computing services to UDs in disaster-stricken areas.
@@
However, several fundamental challenges should be addressed to fully exploit the advantages of UAV-assisted MEC. i) Resource Allocation. Various delay-sensitive and computation-intensive tasks of UDs impose stringent demands on both computing and communication resources, while also exhibiting heterogeneous and time-varying characteristics. However, the limited computing and spectrum resources of UAVs leads to the competition for resources [6, 7]. Consequently, it is challenging to fully utilize the limited resources of UAVs to meet the heterogeneous and stringent demands of UDs. ii) Task Offloading. Due to resource competition, the offloading decision of each UD is influenced not only by its own offloading requirements but also by the offloading decisions of other UDs, resulting in coupled and complex offloading decision-making among UDs [8, 9]. iii) UAV Trajectory Control. While the mobility of UAVs augments the flexibility and elasticity of edge computing services, the limited battery capacity of UAVs inherently restricts the service time [10, 11]. Therefore, the trajectory of UAVs needs to be effectively controlled to balance the service time of UAVs and the QoE of UDs, which presents an additional challenge.
The aforementioned challenges require an efficient joint optimization of resource allocation, task offloading, and UAV trajectory control. However, focusing on just one aspect of these components is insufficient due to the following reasons. On the one hand, due to the limited resources and energy of UAVs, these optimization variables are mutually coupled. For example, the positions of UAVs affect the transmission rate of tasks and the resource allocation influences the completion delay of tasks. Therefore, optimizing task offloading requires simultaneous consideration of resource allocation and the positions of UAVs. On the other hand, these optimization variables collectively determine the QoE of UDs. Therefore, these interconnected optimization variables should be jointly optimized to fully exploit the UAV-assisted MEC system performance, as it can effectively capture the intricate and coupling interactions and trade-offs among various optimization components. To this end, we present a novel online joint optimization approach for task offloading, resource allocation, and UAV trajectory control to maximize the QoE of UDs while adhering to the resource and energy constraints of UAVs. The key contributions of this work are outlined as follows:
@@
  • System Architecture. We propose a hierarchical architecture for multiple-UAV-assisted MEC networks to efficiently coordinate multiple UAVs in providing aerial edge computing services for UDs. Furthermore, we take into account the dynamic mobility and time-varying computational demands of UDs to accurately capture the features of real-world edge computing scenarios within the proposed system.
@@
  • Problem Formulation. We formulate a joint task offloading, resource allocation, and UAV trajectory control optimization problem (JTRTOP), with the aim of maximizing the QoE of UDs under the energy and resource constraints of UAVs. Specifically, the QoE of UDs is theoretically quantified by synthesizing the UD energy consumption and task completion delay. Moreover, the optimization problem is proved to be future-dependent and NP-hard.
@@
  • Algorithm Design. To solve the JTRTOP, we propose a novel online joint task offloading, resource allocation, and UAV trajectory control approach (OJTRTA). Specifically, we first transform the JTRTOP into a per-slot real-time optimization problem (PROP) by using the Lyapunov optimization framework. Then, we propose a two-stage optimization method to jointly optimize task offloading, resource allocation, and UAV trajectory control within the PROP using convex optimization and game theory.
@@
  • Validation. The effectiveness of the proposed OJTRTA is validated through both theoretical analysis and simulation experiments. In particular, the theoretical analysis establishes that the proposed OJTRTA not only satisfies the energy consumption constraints of UAVs, but also has low computational complexity. Additionally, the simulation results affirm the superiority of OJTRTA.
@@
The subsequent sections of this work are structured as follows. In Section II, we review and summarize the related work. Section III details the relevant system models and problem formulation. Section IV describes the Lyapunov-based problem transformation. Section V presents the two-stage optimization method and theoretical analysis. Section VI presents and discusses the simulation results. Finally, this work is concluded in Section VII.
II Related Work
@@
In this section, we provide a comprehensive review of the relevant studies pertaining to UAV-assisted MEC architecture, formulation of joint optimization problems, and online optimization approach. Moreover, we outline the distinctions between this work and the related works in Table I of the supplementary material.
II-A UAV-assisted MEC Architecture
@@
UAV-assisted MEC is a promising paradigm to dynamically expand the computational capabilities of MEC networks and facilitating emergency scenarios. From the perspective of system architecture, existing studies can be broadly classified into single UAV-assisted MEC architecture and multiple-UAV-assisted MEC architecture.
@@
For single UAV-assisted MEC networks, Lin et al. [12] explored the maximization of UAV energy efficiency while simultaneously considering fairness in task offloading within a single UAV-assisted MEC network. Yu et al. [13] presented a single UAV-assisted MEC network, where a UAV is deployed to provide MEC services to Internet of Things (IoT) devices in areas that are inaccessible to the edge cloud due to ground signal blockage or shadowing. Pang et al. [14] proposed a novel mmWave-enabled non-orthogonal multiple access (NOMA) communication scheme, which significantly enhances the energy efficiency of UAV networks. However, a single UAV possesses insufficient coverage range and resources, making it challenging to provide effective services in large-scale areas or areas with a high density of users.
@@
Multiple-UAV-assisted MEC networks, which harness the computational and communication capabilities of multiple UAVs, have been increasingly gaining attention. For example, Li et al. [15] investigated a multiple-UAV-assisted MEC system and devised a robust approach for optimizing computation offloading and trajectory control. Shidrokh et al. [16] conducted a study on resource allocation and task offloading in a multiple-UAV-assisted MEC network, and proposed a two-layer collaborative evolution model to reduce energy consumption and task computation time. Bai et al. [9] proposed a multiple-UAV-enabled edge-cloud computing system that leverages the collaboration between UAVs and remote clouds to deliver exceptional computational capabilities.
@@
II-B Formulation of Joint Optimization Problems
@@ II-B1 Optimization Objectives
Several studies focus on minimizing task completion latency. For instance, Luan et al. [17] presented a joint optimization problem for topology reconstruction and subtask scheduling to minimize the average completion time of subtasks in a UAV-assisted MEC system. Deng et al. [18] investigated the minimization of task service latency in air-ground integrated wireless networks, while considering constraints on learning accuracy and energy consumption.
@@
Several studies focus on minimizing the energy consumption. For example, in the context of a multiple-UAV-assisted two-stage MEC system, Ei et al. [19] aimed to minimize the energy consumption of both mobile devices and UAVs under constraints of task tolerance latency and available resources. Liu et al. [20] conducted a study on a multi-input single-output UAV-assisted MEC network and presented a system energy minimization problem.
@@
There are also the studies that consider the combination of task completion latency and energy consumption as the optimization objective. For example, Tong et al. [21] introduced a UAV-enabled multi-hop cooperative paradigm and formulated an optimization problem to maximize the system utility. The system utility was constructed by jointly considering UAV energy consumption and task completion latency. Hao et al. [22] explored the problem of maximizing the long-term average system gain in a UAV-assisted MEC system, where the system gain was theoretically constructed by combining task completion latency and system energy consumption.
@@
This work differs from the aforementioned studies in the following aspects. First, we considers a more comprehensive set of performance metrics, including task completion latency, UD energy consumption, and UAV energy consumption, which more accurately reflect the practical characteristics of UAV-assisted MEC networks. Furthermore, we combine task completion latency and UD energy consumption into a UD cost function as the optimization objective, while UAV energy consumption is treated as a constraint. This formulation can simultaneously enhance the QoE of UDs and ensure reliable service time for UAVs.
II-B2 Optimization Variables
@@
Researchers have extensively explored various aspects of UAV-assisted MEC systems, with particular emphasis on resource allocation, task offloading, and UAV trajectory control. For example, Chen et al. [23] conducted research on task offloading for IoT nodes and trajectory control for multiple UAVs to reduce UAV energy consumption. Song et al. [24] focused on investigating task offloading and UAV trajectory control problems to improve the performance of UAV-assisted MEC systems. To minimize latency and energy consumption, Pervez et al. [25] jointly optimized task offloading, CPU frequency allocation, transmission power, and UAV trajectory. Seid et al. [26] explored the joint optimization of task offloading, age of information, computation resource allocation, and communication resource allocation to enhance the performance of UAV-enabled Internet of Medical Things networks.
@@
These previous studies have several limitations. First, resource allocation, task offloading, and UAV trajectory control are key factors in enhancing the performance of UAV-assisted MEC networks. However, the aforementioned studies did not comprehensively optimize these factors, which hinders the full exploitation of the advantages of UAV-assisted MEC. Furthermore, these studies consider resource allocation from either the communication or the computation aspects, which may cause performance degradation in practical UAV-assisted MEC systems where both communication and computing resources are insufficient. Motivated by these issues, we jointly optimize task offloading, communication and computing resource allocation, as well as UAV trajectory control.
II-C Online Optimization Approach
@@
To tackle the intricate joint optimization problems of task offloading, resource allocation, and trajectory control for UAV-assisted MEC systems, numerous works have employed offline approaches for system design. These offline approaches formulate solutions under the assumption that the user’s location remains unchanged and the demands of users are either fixed or known in advance [27, 28, 29]. However, in various edge computing scenarios, such as online games and real-time video analysis, computation demands arrive in a stochastic manner, and user locations exhibit dynamic changes. Therefore, it is necessary to develop online approaches to make real-time decisions without knowing future information.
@@
Several studies have also explored online approaches. For example, Zhou et al. [11] developed a two time-scale online approach for caching and task offloading by leveraging the Lyapunov optimization framework. Considering the time-varying computing requirements of user equipment, Wang et al. [30] jointly optimized the user association, resource allocation and trajectory of UAVs with the aim of minimizing energy consumption of all user equipment. To minimize the average power consumption of the system with randomly arriving user tasks, Hoang et al. [31] developed a Lyapunov-guided deep reinforcement learning (DRL) framework. Miao et al. [32] proposed a deep deterministic policy gradient (DDPG)-based algorithm to optimize the computational resources allocation and UAV flight trajectory for UAV-assisted MEC. Xu et al. [7] formulated a long-term optimization problem for the joint optimization of UAV trajectory and computation resource allocation and proposed a trajectory control algorithm based on proximal policy optimization (PPO).
@@
The Lyapunov-based optimization framework and DRL represent two viable methodologies for developing online approaches. While DRL is a powerful technique for training agents to make real-time decisions, it necessitates a substantial amount of sample data to learn optimal strategies and suffers from a lack of interpretability. In contrast, the Lyapunov-based optimization framework does not depend on sample data and provides stable performance guarantees. Therefore, we employ the Lyapunov-based optimization framework to devise our online approach. In a departure from existing research, we propose a novel two-stage optimization method based on game theory and convex optimization within the Lyapunov-based optimization framework. This method demonstrates both low computational complexity and superior performance.
III System Model and Problem Formulation
Figure 1: The hierarchical multiple-UAV-assisted MEC system consists of an aerial control layer, an aerial edge layer and a ground UD layer. Each ground UD can either execute computation tasks locally or offload them to an SUAV or the LUAV for processing. The SUAVs and the LUAV coordinate the allocation of communication and computation resources to receive and execute computation tasks from UDs. Moreover, the LUAV serves as the SDN controller, responsible for collecting system information and making corresponding decisions.